Two planes travel toward each other from cities that are 1100 km apart at rates of 210 km/h and 230 km/h. They started at the same time. In how many hours will they meet?
Answer by amarjeeth123(570)

Let them meet after a time t.
Distance=Speed*time
Distance covered by one plane=210t
Distance covered by the other plane=230t
We have 210t+230t=1100
440t=1100
t=1100/440=110/44=10/4=2.5 hours
They will meet in 2.5 hours.